RHIC Startup: APEX worked through the night. Maintenance Day began at 8:00AM.
Change in the tunes improved the efficiency of both the Blue and Yellow transmission.
Blue Accelerating Cavity #1 went down on Tuesday. A driver in the ring had to be replaced. Beam was lost setting a BLAM alarm.
There was no work done to study the correlation between Blue vertical emittance growth and Air Conditioner cycling. This will be attempted later this week, probably Thursday.
STAR has the FMS trigger implemented as of Monday. It is still being developed. They plan to complete the study of it today, and it will be added to the list of Physics Triggers.
PHENIX has collected more than half of the total luminosity they require for d-Au running. They have reversed the polarity of their central magnet. This will require retuning of the orbit, so a 6 bunch ramp will be brought up first to make these corrections. In the past this required a change in the orbit and coupling. This year feedback on the ramp complicates matters. Last year we never recovered the same rates as before the field reversal. There is also a question of whether or not the electron cloud will create vacuum problems in PHENIX. we have not seen bad vacuum effects in PHENIX yet this year.
Scheduling Physicist: Mike Sivertz | Last updated: |
Email: sivertz@bnl.gov | |
Office Phone: 631-344-6102 | |
Digital Pager: 7528 or 4300 | |
Home Phone: 631-878-9020 |